A house with a mains supply at 250 V has three 3 KW electric heaters and ten 60 W lamps. The power and the lighting circuits are separate, and each has its own main fuse . What current passes through each of the fuses when all the heaters and lamps are in use?
1
Expert's answer
2020-04-07T09:19:15-0400
"I_H=\frac{P_H}{V}=\frac{3(3000)}{250}=36\ A"

"I_L=\frac{P_L}{V}=\frac{10(60)}{250}=2.4\ A"